OpenStudy (anonymous):

19.At a dairy it takes 90 seconds to fill 30 one-gallon jugs of milk. How many minutes does it take to fill 90 jugs of milk? A. 1 B. 3 C. 3.5 D. 4 E. 4.5

hero (hero):

Hint: Figure out how many seconds it takes to fill 90 jugs, then convert seconds into minutes.

OpenStudy (anonymous):

thats the question.

OpenStudy (jaja):

answer is E.

OpenStudy (jaja):

now go help me

OpenStudy (anonymous):

@jaja how explain please.

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

@KingGeorge help here please.

hero (hero):

Second hint: Set up a proportion: \(\frac{90 \space\text{seconds}}{x \space\text{seconds}} = \frac{30 \space\text{jugs}}{90 \space\text{jugs}}\) Solve for x, then convert to minutes

OpenStudy (anonymous):

4.5

OpenStudy (anonymous):

90/30 *90sec=90/30 *1.5min=4.5 min

OpenStudy (anonymous):

Time (Sec) : Milk Jugs = Time(Sec) : Milk Jugs 90 : 30 = x : 90 So: \[\frac{90}{30} = \frac{x}{90}\] \[\frac{x}{90} = 3\] \[x = 270\; (Seconds)\] So to convert it into minutes divide it by 60: \[270 \; Seconds = \frac{270}{60} \implies \color{blue}{4.5 \; \; minutes.. }\]